OpenStudy (anonymous):

x^2-5x-36 Use Quadratic Formula

OpenStudy (anonymous):

\[{x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]

OpenStudy (anonymous):

its not the answer it has to come out to a answer

OpenStudy (anonymous):

a Number

OpenStudy (anonymous):

ahh.. I thought you don't know the formula :F lol...

OpenStudy (anonymous):

lol yea

OpenStudy (tommo_lcfc):

x^2 - 5x - 36. x = 5 + or - sqrt(25-4(36))/ 2 5/2 +/- sqrt(119)/2 i This has complex roots so you can't get a real number. Hope this helps :)

OpenStudy (anonymous):

you have: \[\LARGE x^2-5x-36=0\] now substitute: \[\LARGE {x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] \[\large {x_{1/2}} = \frac{{ - (-5) \pm \sqrt {{25} - 4(-36)} }}{{2}}\] can you do it now ? :)

OpenStudy (anonymous):

18/2 and -8/2

OpenStudy (tommo_lcfc):

Sorry, I forgot it was -36. Follow @Kreshnik. I don't know how to do it like that!

OpenStudy (anonymous):

@hoggard1020 need more help? :)

OpenStudy (anonymous):

hoggard, you ok now?

OpenStudy (tommo_lcfc):

@thomasj, this simplifies down to 9 and -4!

OpenStudy (anonymous):

yes its diffucult

OpenStudy (anonymous):

ahh don't worry, it's pretty easy, you'll see ;) \[\LARGE {x_{1/2}} = \frac{{ - (-5) \pm \sqrt {{25} - 4(-36)} }}{{2}}\] \[\LARGE {x_{1/2}} = \frac{{ 5 \pm \sqrt {{25} +144} }}{{2}}\] \[\LARGE {x_{1/2}} = \frac{{ 5 \pm \sqrt {{169}} }}{{2}}\] \[\LARGE {x_{1/2}} = \frac{{ 5 \pm 13 }}{{2}}\] NOW YOU HAVE: \[\LARGE {x_{1}} = \frac{{ 5 - 13 }}{{2}}\] AND \[\LARGE {x_{2}} = \frac{{ 5 + 13 }}{{2}}\] GO FOR IT...

OpenStudy (anonymous):

Thanks =)

OpenStudy (anonymous):

in the end you should have @tommo_lcfc 's result :)

OpenStudy (anonymous):

OUR pleasure xD

OpenStudy (tommo_lcfc):

I would not use the quadratic formula normally unless the equation could not be factorised. In this case, it can be as 9 and -4 are suitable factors of 36 summing to 5 and producing -36 when multiplied. If the question says to use the quadratic formula as it did, you obviously should use it but if not, factorise if you can because that is easier and less time consuming.

OpenStudy (tommo_lcfc):

in my opinion

OpenStudy (anonymous):

@tommo_lcfc it saves time when you have simple quadratic formula... how about this: can you factorize this out? \[\LARGE x^2-14x-72=0\]

OpenStudy (tommo_lcfc):

OK, you win this round.